I need to find the exponential smoothing.
0.3 0.8 INOV Loct 131-175 [130.074 Dec 129.47 Jan 128.577 Feb 128,06 march 125. 136 1
Answer #1

Ans) Formula is, ft+1 =αdt + (1-α)ft

where,

dt =actual demand and

ft= previous forecast

For ,α1 =0.3

november=0.3*131.175 +(1-0.3)131.175 = 131.175

december = 0.3 * 130.074 + 0.7 * 131.175 = 130.844

january = 0.3 * 129.47 + 0.7 * 130.844 = 130.431

February = 0.3 * 128.577 + 0.7 * 130.431= 129.874

March = 0.3 * 128.06 + 0.7 * 129.874 = 129.329

October = 0.3 * 125.136+ 0.7 * 129.329 = 127.936

For, α2 =0.8

november=0.8*131.175 +(1-0.8)131.175 = 131.175

december = 0.8 * 130.074 + 0.2 * 131.175 = 130.294

january = 0..8* 129.47 + 0.2 * 130.294 = 129.634

February = 0.8 * 128.577 + 0.2 * 129.634= 128.788

March = 0.8 * 128.06 + 0.2 * 128.788 = 128.205

October = 0.8 * 125.136+ 0.2 * 128.205 = 125.749

Forecast for 0.3 Forecast for 0.8
October 131.175 127.936 125.749
November 130.074 131.175 131.175
December 129.47 130.844 130.294
January 128.577 130.431 129.634
February 128.06 129.874 128.788
March 125.136 129.329 128.205

Do give positive rating if this answer is helped.